OpenStudy (volleyballlover55):

Convert 0.0001 to a power of 10. 10^−4 10^4 10^−3 10^3

OpenStudy (volleyballlover55):

A?

OpenStudy (shamim):

0.1=?

OpenStudy (shamim):

0.1=1/10=1*10^-1 Right?

OpenStudy (volleyballlover55):

right

OpenStudy (volleyballlover55):

so its A

OpenStudy (shamim):

Result is not all

OpenStudy (volleyballlover55):

what?

OpenStudy (shamim):

We r going to dscuss nd tr to learn

OpenStudy (shamim):

Now tell me 0.001=?

OpenStudy (volleyballlover55):

ok but i umderstand.. i know the answer is A

OpenStudy (shamim):

Then plz tell me 0.0001=?

OpenStudy (volleyballlover55):

10^-4

OpenStudy (shamim):

0.00001=?

OpenStudy (shamim):

Yes u r correct!!!

OpenStudy (volleyballlover55):

thanks next!

OpenStudy (volleyballlover55):

Convert 5.4 ⋅ 10^3 to standard form. 0.0054 0.00054 5,400 54,000

OpenStudy (volleyballlover55):

C?

OpenStudy (shamim):

U r most welcome!!!

OpenStudy (volleyballlover55):

please help me with more.. i know all the answers to the questions i going to ask you.. i just need to make sure they are correct :)

OpenStudy (shamim):

Plz post ur new problem again

OpenStudy (volleyballlover55):

Convert 5.4 ⋅ 10^3 to standard form. 0.0054 0.00054 5,400 54,000